OpenStudy (anonymous):

Solve for x: |x| − 8 = −5

jimthompson5910 (jim_thompson5910):

Any ideas on the first step?

OpenStudy (anonymous):

add both by 8

jimthompson5910 (jim_thompson5910):

good so you get |x| − 8+8 = −5+8 which becomes |x| = 3

OpenStudy (asnaseer):

@TuringTest has already explained to you how to solve this exact question here: http://openstudy.com/updates/4ff8c4d0e4b058f8b76339f5

jimthompson5910 (jim_thompson5910):

So what's next?

OpenStudy (anonymous):

dnt kno jit

jimthompson5910 (jim_thompson5910):

Use the idea that if |x| = k for some positive number k, then x = k or x = -k